OpenStudy (anonymous):

When using the Pythagorean Theorem to find the hypotenuse of the triangle, do you have to square your answer you got after you squared and added all the numbers?

OpenStudy (yamyam70):

No, you have to take it , like this.. c^2 = a ^2 + b^2 ,, then...

OpenStudy (yamyam70):

\[c = \sqrt{a^2 + b^2}\]

OpenStudy (anonymous):

Yep got it I was guessing the formula but I get it thanks!